OpenStudy (anonymous):

Solve the equation. 1 over 5x + 15 minus 3 over x + 3 = 5 a) x = negative 89 over 25 b) x = 89 over 25 c) x = negative 61 over 25 d) x = 61 over 25

OpenStudy (anonymous):

\[\frac{ 1 }{ 5x+15 }-\frac{ 3 }{ x+3 } = 5\] is this your question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we start by finding a common denominator, this is easy because all we have to do is multiply the numerator and denominator of 3/(x+3) by 5, so we then get : 1/(5x+15) - 15/(5x+15) = 5 now that we have a common denominator, we can subtract to get -14/(5x+15) = 5 multiply by the denominator to remove x from the denominator, we get: -14 = 25x +75 subtract 75 from both sides: 89 = 25x divide by 25 to get 89/25 = x

OpenStudy (anonymous):

thank you thank you thank you very much!!!!!!!!!!!!!

OpenStudy (anonymous):

you're welcome!